Multiply (9+ the square root of 10) by (9- the square root of 10). Simplify and type an exact answer using radicals as needed.
Answer by algebrahouse.com(1659)

"Multiply (9+ the square root of 10) by (9- the square root of 10). Simplify and type an exact answer using radicals as needed."

Use the foil method when multiplying two binomials.
(9 + √10)(9 - √10)
= 9(9) + 9(-√10) + √10(9) + √10(-√10) {used foil method}
= 81 - 9√10 + 9√10 - √100 {multiplied through}
= 81 - √100 {combined like terms}
= 81 - 10 {evaluated square root of 100 to be 10}
= 71

www.algebrahouse.com